40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 287 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  3170   Wed Jul 7 17:18:57 2010 AlbertoConfigurationElectronicsStochmon and LSC AM Stabilizer Decomissioned
Today I disconnected and removed the Stochmon box from the 1Y2 rack.
I also removed the amplifiers that were sitting on the PSL table, next to the RF AM PD, that were connected to the Stochmon. I pulled back the RG cable and the power cables that went from the PSL table to the 1Y1 and 1Y2 racks.
The power cable, all rolled up, is now sitting on the floor, inside the 1Y1 rack and one of its end is still connected to the power of the rack. We'd like to turn off the entire rack in order to safely remove it. But since the laser driver is there too, we should do it the first time we have to turn off the rack for some other reason.

I also removed two of the AM stabilizers from the 1Y2 rack. The other one, which is currently running th MC modulations, is still in the rack, and there it is going to remain together with its distribution box.

I stored both AM stabilizers and the Stochmon box inside the RF cabinet down the East arm.

  3169   Wed Jul 7 17:05:30 2010 steve, ranaUpdateSAFETYSummary of Crane Damage/Malfunction

The 1 Ton yellow crane support beam jammed up at Friday morning, June 25.

 

The 40m vertex crane has a folding I-beam support to reach targeted areas. The rotating I-beam is 8 ft long. The folding extension arm gives you another 4 ft.

The 12 ft full reach can be achieved by a straightening of the 4 ft piece. There is a spring loaded latch on the top of the I-beam that locks down when the two I-beams align.

This lock joins the two beams into one rigid support beam for the jib trolley to travel.  The position of this latch is visible when standing below, albeit not very well.

To be safe it is essential that this latch is locked down fully before a load is put on the crane.

 

We were preparing to pump down the 40m vacuum system on Friday morning. The straight alignment of the 8 and 4 ft piece made us believe that

the support beams were locked. In reality, the latch was not locked down. The jib trolley was driven to the end of the 12 ft I-beam. The 200 lbs ITM-east door was lifted

when the 4 ft section folded 50 degrees around the pivot point. This load of door + jib-trolley + 4 ft I-beam made the support beam sag about 6 inches

The door was removed from the jib hoist with the blue Genie-lift. The sagging was reduced to ~3".

 

The Genie-lift platform was raised to support the sagging crane jib-trolley. The lab was closed off to ensure safety and experts were called in for consultation. It was decided to bring in professional riggers.

 

Halbert Brothers, Inc. rigging contractor came to the lab Tuesday morning to fix the crane. The job was to unload the I-beam with safety support below. They did a very good job.

 

The static deformation of I-beams sprung back to normal position. There are some deformation of the I-beam ~2 mm where the beams were jammed under load.

It is not clear if this is a new deformation or if the crane sections have always been mis-aligned by a couple of mm.

 

The crane was tested with 450 lbs load at 12 ft horizontal travel position. The folding of I-beams were repeatedly tested for safe operation. Its a 1 ton crane, but we tested it with 450 lbs because that's what we had on hand.

 

We're working on the safety upgrade of this lift to prevent similar accident from happening.

Pictures below:

Atm 1)   load testing 2007

Atm 2)   jammed-sagging under ~400 lbs, horizontal          

Atm.3)   jammed-folded 50 degrees, vertical     

Atm.4)   static deformation of I-beams

Atm.5)   unloading in progress with the help of two A-frames          

Atm.6)   it is unloaded

Atm.7-8) load testing        

Atm.9)   latch locked down for safe operation            

Atm.9)   zoom in of the crane sections misalignment    

Attachment 1: DSC_0026_00.JPG
DSC_0026_00.JPG
Attachment 2: P1060408.JPG
P1060408.JPG
Attachment 3: P1060415.JPG
P1060415.JPG
Attachment 4: P1060413.JPG
P1060413.JPG
Attachment 5: P1060421.JPG
P1060421.JPG
Attachment 6: P1060423.JPG
P1060423.JPG
Attachment 7: P1060441.JPG
P1060441.JPG
Attachment 8: P1060436.JPG
P1060436.JPG
Attachment 9: P1060425.JPG
P1060425.JPG
Attachment 10: P1060432.JPG
P1060432.JPG
  3168   Wed Jul 7 12:45:00 2010 nancyUpdate Weekly Update

Wednesday after the meeting - Started report, learnt mode cleaner locking from Kiwamu and Rana, saw how to move optics on the tables with Rana and kiwamu.

Thursday - Made the report

Tuesday - report.

Today - am trying locking the MC with kiwamu's help to see the WFS signals and also to start characterizing the QPD.

  3167   Wed Jul 7 12:17:36 2010 RazibUpdatePhase CameraWeekly update

I have completed the following tasks:

1. Find a simplified calibration of the exposure time for the current He-Ne setup. Basically, I triggered the camera to take 20 snapshots with a 20 Hz driving signal at different exposure time beginning from 100 us (microsecond) upto 4000 us with an increment of 200 us.

    Result: The current power allows the camera to have an exposure time of ~500 us before the DC level begans to saturate.

2. Aidan and I did some alignment and connected the AOM and corrected the driving frequency of its PZT to 40 Mhz(which apparently was disconnected which in turn gets the credit of NO beat signal for me until Tuesday 07/06/2010 5:30 PST) .

    Result: I got the beat signal of 1 Hz and 5 Hz. Image follows (the colormap shows the power in arbitrary units).

3. Finished writing my Progress Report 1 .

DC_1Hz_beat_sig.jpgDC_5Hz_beat_sig.jpg

Attachment 1: DC_1Hz_beat_sig.jpg
DC_1Hz_beat_sig.jpg
  3166   Wed Jul 7 11:35:59 2010 GopalUpdateWIKI-40M Update6.30.10 - 7.7.10 Weekly Update

Summary of this Week's Activities:

6/30: 2nd and 3rd drafts of Progress Report

7/1: 4th draft and final drafts of Progress Report; submitted to SFP

7/5: Began working through busbar COMSOL example

7/6: LIGO meeting and lecture; meeting with Koji and Steve to find drawing of stacks; read through Giaime's thesis, Chapter 2 as well as two other relevant papers.

7/7: Continued working on busbar in COMSOL; should finish this as well as get good headway on stack design by the end of the day.

  3165   Wed Jul 7 11:23:08 2010 Sharmila, KatharineUpdateWIKI-40M Update 

Weekly Update:

Last Weds-Thurs, we wrote and edited our progress reports.

Tuesday (and Weds morning):  Continued circuit analysis of Haixing's circuit and plotting transfer functions (almost have one for entire circuit).  Hooked up OSEM and circuit to power supply, but the LED didn't appear to light up in IR.  Now we are going to hook the OSEM directly to the power supply, sans circuit, to see if the problem is with the circuit or OSEM.

  3164   Wed Jul 7 10:42:29 2010 KojiSummaryPSLpower spectral density from RefCav transmitted beam

How do you calibrate this to Hz/rtHz?

Quote:

I measured the RC transmitted light signals here at the 40m. I made all connections through the PSL patch panel. No optics/PD were touched.

I measured the spectral density of the signal of the transmitted beam behind RefCav in both time and frequency domain.

This will be compared with the result from PSL lab later, so I can see how stable the signal should be.

We re-aligned the beam into the cavity (the DC level increased from 2 V to 3.83V)

and the reflected beam to the center of the RFPD.

 

 

  3163   Wed Jul 7 00:15:29 2010 tara,RanaSummaryPSLpower spectral density from RefCav transmitted beam

I measured the RC transmitted light signals here at the 40m. I made all connections through the PSL patch panel.

Other than two steering mirrors in front of the periscope, and the steering mirror for the RFPD which were used to steer

the beam into the cavity and the RFPD respectively, no optics are adjusted.

We re-aligned the beam into the cavity (the DC level increased from 2 V to 3.83V) (Fig2) (We could not recover the power back to what it was 90 days ago)

and the reflected beam to the center of the RFPD.

 

I measured the spectral density of the signal of the transmitted beam behind RefCav in both time and frequency domain.

This will be compared with the result from PSL lab later, so I can see how stable the signal should be.

I did not convert Vrms/rtHz to Hz/rtHz because I only look at the relative intensity of the transmitted beam which will be compared to the setup at PSL lab. 

 

 

 We care about this power fluctuation because we plan to measure

 photo refractive noise on the cavity's mirros

(this is the noise caused by dn/dT in the coatings and the substrate,

the absorption from fluctuating power on the coating/mirror changes

the temperature which eventually changes the effective length of the cavity as seen by the laser.)

      

      The plan is to modulate the power of the beam going into the cavity,

the absorption from ac part will induce frequency noise which we want to see.

Since the transmitted power of the cavity is proportional to the power inside the cavity.

 Fluctuations  from other factors, for example, gain setting,  will limit our measurement. 

That's why we are concerned about the stability of the transmitted beam and made this measurement.


 

Attachment 1: RIN_rftrans.png
RIN_rftrans.png
Attachment 2: tara.png
tara.png
  3162   Tue Jul 6 17:38:27 2010 JenneUpdateSUSTip Tilt Progress!

[Jenne, Kyung Ha]

We successfully suspended the 4 eddy current dampers for the first Tip Tilt.  We had some lessons learned, including how to carefully get an allen wrench in between the dampers to do up some of the screws, and how to be careful not to bend the wire while tightening the screws.  More tomorrow...

  3161   Tue Jul 6 17:29:04 2010 ChipUpdatePEMThe Ranger is mine!
  3160   Tue Jul 6 17:07:56 2010 josephbUpdateCDSc1sus status

I talked to Alex, and he explained the steps necessary to get the real time linux kernel installed.  It basically went like copy the files from c1iscex (the one he installed last month) in the directory /opt/rtlidk-2.2 to the c1sus locally.  Then go into rtlinux_kernel_2_6, and run make and make install (or something like that - need to look at the make file).  Then edit the grub loader file to look like the one on c1iscex (located at /boot/grub/menu.lst).

This will then hopefully let us try out the RCG code on c1sus and see if it works.

  3159   Tue Jul 6 17:05:30 2010 Megan and JoeUpdateComputersc1iovme reboot

We rebooted c1iovme because the lines stopped responding to inputs on C1:I00-MC_DRUM1. This fixed the problem.

  3158   Tue Jul 6 11:57:06 2010 josephbUpdateCDSDaily Downs Update

I went to talk to Rolf and Jay this morning.  I asked Rolf if a chassis was available, so he went over and disconnected one of his test stand chassis and gave it to me.  It comes with a Contect DIO-1616L-PE Isolated Digital IO board and an OSS-MAX-EXP-ELB-C, which is a host interface board.  The OSS board means it has to go into the south end station.  There's a very short maximum cable length associated with that style, and the LSC and IOO chassis will be further than that from their computers (we have dolphin connectors on optical fiber for those connections).

I also asked Jay for another 4 port 37 d-sub ADC blue and gold adapter box, and he gave me the pieces.  While over there, I took 2 flat back panels and punched them with approriate holes for the scsi connectors that I need to put in them.  I still need to drill 4 holes in two chassis to mount the boards, and then a bit of screwing.  Shouldn't take more than an hour to put them both together.  At that point, we should have all the adapter boxes necessary for the base design.  We still need some stuff for the green locking, as noted on Friday.

Major hardware still needed:

2 Dolphin style IO chassis

1 computer for south end front end

  3157   Fri Jul 2 11:33:15 2010 josephbUpdateCDSc1sus needs real time linux to be setup on it

I connected a monitor and keyboard to the new c1sus machine and discovered its not running RTL linux.  I changed the root password to the usual, however, without help from Alex I don't know where to get the right version or how to install it, since it doesn't seem to have an obvious CD rom drive or the like.  Hopefully Tuesday I can get Alex to come over and help with the setup of it, and the other 1-2 IO chassis.

  3156   Fri Jul 2 11:06:38 2010 josephb, kiwamuUpdateCDSCDS and Green locking thoughts

Kiwamu and I went through and looked at the spare channels available near the PSL table and at the ends.

First, I noticed I need another 4 DB37 ADC adapter box, since there's 3 Pentek ADCs there, which I don't think Jay realized.

PSL Green Locking

Anyways, in the IOO chassis that will put in, for the ADC we have a spare 8 channels which comes in the DB37 format.  So one option, is build a 8 BNC converter, that plugs into that box.

The other option, is build 4-pin Lemo connectors and go in through the Sander box which currently goes to the 110B ADC, which has some spare channels.

For DAC at the PSL, the IOO chassis will have 8 spare channel DAC channels since there's only 1 Pentek DAC.  This would be in a IDC40 cable format, since thats what the blue DAC adapter box takes.  A 8 channel DAC box to 40 pin IDC would need to be built.

 

End Green Locking

The ends have 8 spare DAC channels, again 40 pin IDC cable.   A box similar to the 8 channel DAC box for the PSL would need to be built.

The ends also have spare 4-pin Lemo capacity.  It looked like there were 10 channels or so still unused.  So lemo connections would need to be made.  There doesn't appear to be any spare 37 DB connectors on the adapter box available, so lemo via the Sander box is the only way.

 

Notes

Joe needs to provide Kiwamu with cabling pin outs.

If Kiwamu makes a couple spares of the 8 BNC to 37DB connector boards, there's a spare 37DB ADC input in the SUS machine we could use up, providing 8 more channels for test use.

  3155   Thu Jul 1 17:05:36 2010 josephb, kiwamu, steveUpdateCDSNew SUS Chassis in rack

Thanks to Steve's work on some L brackets, and Kiwamu's lifting help, we now have a new SUS IO chassis in the new 1X4 rack (formerly the 1Y4 rack), just below the new SUS and LSC computers.  I have decided to call the sus machine, c1sus, and give it IP address 192.168.113.85.  We also put in a host interface adapter, OSS-HIB2-PE1x4-1x4 Re-driver HIB, which connects the computer to the IO chassis.

The IP was added to the linux1 name server.  However, the computer itself has not been configured yet.  I'm hoping to come in for an hour or two tomorrow and get the computer hooked up to a monitor and keyboard and get its network connection working, mount /cvs/cds and get some basic RCG code running.

We also ran ethernet cables for the SUS machine to the router in 1X6 (formerly 1Y6) as well as a cable for megatron from 1X3 (formerly 1Y3) to the router, in anticipation of that move next week.

During the day, I realized we needed 2 more ADCs, one of which I got from Jay immediately.  This is for two 110Bs and 4 Pentek ADCs.  However, there's a 3rd 110B connected to c0dcu1 which goes to a BNC patch panel.  Original Jay thought we would merge that into 4 pin lemo style into the 2nd 110B associated with the sus front ends.  We've decided to get a another ADC and adapter.  That will have to be ordered, and generally take 6-8 weeks.  However, it may be possible to "borrow" one from another project until that comes in to "replace" it.  This will leave us with our BNC patch panel and not force me to convert over 20 cables.

I also discovered we need one Contec DIO-1616L-PE Isolated Digital IO board for each Chassis, which I wasn't completely aware of.  This is used to control the ADCs/DACs adapter boards in the chassis.  It means we need still need to put a Binary Output board in the c1iscex chassis.  Hopefully the chassis as they come in come from Downs continue to come with the Contec DIO-1616L-PE boards (they have so far).

The current loadout of the SUS chassis is as follows:

Treton Board

Far left slot, when looking from the front has the OSS-MAX-EXP-ELB-C board, used to communicate with the c1sus computer.

Slot 1 ADC PMC66-16AI6455A-64-50M

Slot 2 DAC PMC66-16AO16-16-F0-OF 

Slot 3-6 BO Contec DIO-32L-PE Isolated Digital Output board

Slot 7 ADC PMC66-16AI6455A-64-50M

Slot 8-9 DAC PMC66-16AO16-16-F0-OF 

Slot 10-11 ADC PMC66-16AI6455A-64-50M

Slot 12 Contect DIO-1616L-PE Isolated Digital IO board

Back  Board

Slot 1 ADC adapter D0902006

Slot 2 DAC adapter D0902496-v1

Slot 7 ADC adapter D0902006

Slot 8-9 DAC adapter D0902496-v1

Slot 10-11 ADC adapter D0902006

 

  3154   Thu Jul 1 14:28:39 2010 JenneUpdatePEMMC1 Accelerometers in place

Kevin sent me an email with top secret info on where one of the other accelerometer cubes was hiding (it was with his shaker setup on the south side of the SP table), so I took it and put the 3 MC1 accelerometers in their 3-axis configuration. 

Also, I changed the orientation of both sets of 3 axis accelerometers to reflect a Right Handed configuration, to go along with the new and improved IFO configuration.  Previously (including last night), the MC2 accelerometers were together in a Left Handed configuration.

  3153   Thu Jul 1 14:03:40 2010 josephbUpdatePEMTemporary disconnect of some PEM channels for ~20 minutes

In order to identify the output adapter of the BNC patch panel used for about 20 PEM channels, I had to disconnect its power and remove the back panel.  Channels coming into that panel (seismometers and so forth) was out from 1:36 to 1:56 pm.

I did a quick check of some of the channels and it looks like its working again after putting it all back together.

  3152   Thu Jul 1 10:26:34 2010 steveConfigurationVACpd68 at vacuum normal state

Slow pump down _pd68 has reached the vacuum normal state.  CC4_Rga region is pumped now. The RGA is still off.

Precondition to this pump down:                                                                                                                                                                                                                                                           129 days at atm,                                                                                                                                                                                                                                                                                         ITMs replaced. MMT, oplev and other components were removed from BSC, ITMCs. New MMT mirrors are in.                                                                                                             IOO_access_connector was out. The end chambers were not opened.

 

Attachment 1: pd68slow2d.jpg
pd68slow2d.jpg
Attachment 2: pem134d.jpg
pem134d.jpg
  3151   Wed Jun 30 23:03:46 2010 ranaConfigurationIOOPower into MC restored to max

Kiwamu, Nancy, and I restored the power into the MC today:

  1. Changed the 2" dia. mirror ahead of the MC REFL RFPD back to the old R=10% mirror.
  2. Since the MC axis has changed, we had to redo the alignment of the optics in that area. Nearly all optics had to move by 1-2 cm.
  3. 2 of the main mounts there had the wrong handedness (e.g. the U100-A-LH instead of RH). We rotated them to some level of reasonableness.
  4. Tuned the penultimate waveplate on the PSL (ahead of the PBS) to maximize the transmission to the MC and to minimize the power in the PBS reject beam.
  5. MC_REFL DC  =1.8 V.
  6. Beams aligned on WFS.
  7. MC mirrors alignment tweaked to maximize transmission. IN the morning we will check the whole A2L centering again. If its OK, fine. Otherwise, we'll restore the bias values and align the PSL beam to the MC via the persicope.
  8. waveplates and PBS in the PSL were NOT removed.
  9. MC TRANS camera and QPD have to be recentered after we are happy with the MC axis.
  10. MC REFL camera has to be restored.
  11. WFS measurements will commence after the SURF reports are submitted.

We found many dis-assembled Allen Key sets. Do not do this! Return tools to their proper places or else you are just wasting everyone's time!

 

  3150   Wed Jun 30 18:16:55 2010 steveUpdateSAFETYSURF safety training 2010

40m SURFs Razib Obaid, Nancy Aggarwal, Unknown Bearded SMURF, Megan Daily, Gopal Nataraj, Katharine Larson and Sharmila Dhevi received 40m specific safety training on June 23, 2010.

Attachment 1: DSC_1902.JPG
DSC_1902.JPG
  3149   Wed Jun 30 17:30:54 2010 JenneUpdatePEMSeismometers moved for MC Wiener work

Now that the MC is back up and running, I put the Guralp seismometers at the ends of the mode cleaner.  Gur1 is near MC2, and Gur2 is near MC1 (yes, it seems backwards....that's how the cable lengths work).  Also, the set of 3 MC2 accelerometers are in place under MC2.  I can't find the black cube for the other set of accelerometers, so there aren't any around MC1/3. 

  3148   Wed Jun 30 15:24:04 2010 steve,kiwamuUpdateVACslow pumpdown copmlete

Quote:

Untitled.png

The pumpdown started at 4 PM (2300 UTC). At 10 PM, we (Jenne, Jan, and I) opened up the RV1 valve to full open. That's the second inflection point in the plot.

 Atm 2 is showing the butterfly valve that closes down down the orifice at higher pressure to slow down the pumping speed.

 See elog entry #2573

 

Attachment 1: slowpd.jpg
slowpd.jpg
Attachment 2: butterfly.JPG
butterfly.JPG
  3147   Wed Jun 30 14:38:04 2010 JenneUpdatePEMGuralp terminator is switched

I moved the Guralp box's input terminator from Gur1 to Gur2 a minute or so ago to check the other channels.

  3146   Wed Jun 30 12:20:49 2010 RazibUpdatePhase CameraWeekly update

This week I have completed following tasks:

1. Worked out the analytical expressions for the amount of power of the DC and oscillatory part going into the camera.

2. Realigned the He-Ne PhaseCam setup as we had to replace the first steering mirror after the laser with a silvered mirror ( one without a dielectric coating for 1064 nm).

3. Gone through the code written by a previous surfer (Zach Cummings).

4. Read the paper 'Real-time phase-front detector for heterodyne interferometers'- F. Cervantes et. el. where they talk about constructing a phase detector for LISA pathfinder mission. One interesting fact I found was that, they used InGaAs chip for their CCD Cam which has a amazing QE of 80% @ 1064 nm. Unfortunately, the one we are using (Micro MT9V022 CMOS) has only ~5% QE for 1064 nm and 50% for 633 nm. One top of it MT9V022 has a built-in infra-red filter infront of it to make it more insenstive to 1064. In such limitations, we may have to find a work-around for this issue. Any idea?

5. Read about the EOM and AOM and their vibrating (!) way to add on and alter the incident light on them. (Source: Intro to Optical Electronics-Yariv)

 

One task that we couldn't accomplish even though I planned on doing is:

1. Move,if possible, to the Nd:YAG setup.

 

Task for this week:

1. Produce breathtaking calibration of the camera at He-Ne setup.

2. Read 'Fringe Analysis'-Y.Surrel and 'Phase Lock Technique'-Gardner.

3. Modify the phasecam code.

4. Produce an alternate triggerbox using diodes instead of Op-Amp as op-amp is suspected to fail at some point driving the camera due to impedance mismatch.

5. Answer Koji's question at Aidan's ELOG .

  3145   Wed Jun 30 12:13:39 2010 Sharmila, KatharineUpdateWIKI-40M Updateweekly update

Weekly Project Update:

We are studying Haixing's circuit diagram for the quadrant maglev control circuit.  We have analyzed several of the sub-circuits and plotted transfer functions for these in MatLab.  To check the circuit, we will compare the calculated transfer functions with those obtained from the HP control systems analyzer.

To learn how to use the control systems analyzer, we are reading App Note 243 as well as an online manual (477 pages in the first volume).  We are beginning with a simple test circuit, and are comparing its measured frequencyresponse with calculated transfer functions.  We currently have obtained a response graph beginning at 100 Hz (which we have not yet figured out how to print), and we are planning to investigate behavior at lower frequencies.

We also have been continuing our reading on control systems after a failed attempt at magnetic levitation. 

  3144   Wed Jun 30 12:01:20 2010 josephbUpdateCDSSUS IO Chassis

I spent this morning populating the SUS IO Chassis and getting it ready for installation into the 1Y4 rack.  I discovered I'm lacking the internal cables to connect the ADC/DAC boards to the AdL adapter boards that also go in the chassis (D0902006 and D0902496).  I'm not sure where Alex got the cables he used for the end IO chassis he had put in.  I'll be going to Downs after the meeting today either to get cables or parts for cables, and get the SUS chassis wired up fully.

I'd also like to confirm with Alex that the OSS-MAX-EXP-ELB-C board that goes in the IO chassis matches the host interface board that goes in the computer (OSS-HIB2-PE1x4-1x4 Re-driver HIB, since we spent half a day the last time we installed an IO chassis determining that one of the pair was bad or didn't match.

 

The SUS chassis has been populated in the following way:

Treton Board

Slot 1 ADC PMC66-16AI6455A-64-50M

Slot 2 DAC PMC66-16AO16-16-F0-OF 

Slot 3-6 BO Contec DIO-1616L-PE Isolated Digital IO board

Slot 7 ADC PMC66-16AI6455A-64-50M

Slot 8-9 DAC PMC66-16AO16-16-F0-OF 

Back  Board

Slot 1 ADC adapter D0902006

Slot 2 DAC adapter D0902496-v1

Slot 7 ADC adapter D0902006

Slot 8-9 DAC adapter D0902496-v1

  3143   Wed Jun 30 11:39:20 2010 nancyUpdateWIKI-40M UpdateWeekly Update

Wednesday Morning E-log :

 

Most of the time through this week, i was working towards making the simulink model work.

It involved learning simulink functions better, and also improving on the knowledge of control theory in general, and control theory of our system.

1. Thusrday : found tfs for the feedback loop. and tried many different filters and gains to stabilize the system (using the transient response of the system). - not through

2. Friday : decided to use error response and nullify the steady state error instead of looking at convergence of output. tried many other filter functions for that.

Rana then showed me his files for WFS.

3. Sunday - played with rana's files, learnt how to club simluink with matlab, and also about how to plot tfs using bode plots in matlab.

4. Monday : Read about state-space models, and also how to linearize in matlab. done with the latter, but the former still needs deeper understanding.

read ray-optics theory to calculate the geometric sensing matrix.

It first requires to calculate the eigen mode of the cavity with tilted mirrors. this eigen mode is needed to be found out using ray-optics transfer matrices for the optics involved  . figured out  matrices for the tilted plane mirrors, and am working on computing the same for MC2.

5. Tuesday : went to Universal Studios , Hollywood :P

6. Wednesday (today) : Writing the report to be submitted to SFP.

  3142   Wed Jun 30 11:35:06 2010 Gopal UpdateGeneral6.23.10 - 6.30.10 Weekly Update

Summary of this Week's Activities:

6/23: LIGO Safety Tour; Simulink Controls Tutorial

6/24: Simulink Diagram for Feedback Loop; Constructed Pendulum Transfer Function; Discussion with Dr. Weinstein

6/25: Prepare for pump-down of vacuum chamber; crane broken due to locking failure; worked through COMSOL tutorials

6/28: Ran through Python Tutorials; Began learning about Terminal

6/29: Wrote Progress Report 1 First Draft

6/30: Began editing Progress Report 1

  3141   Wed Jun 30 00:45:26 2010 ranaUpdateVACslow pumpdown stopped for the night

As per Steve's instructions, at 12:43 AM, I used the following steps to stop the pumpdown until the morning:

  1. Close RV1 using the 'steering-wheel' wrench.
  2. Close V3.
  3. Turn OFF RP1.
  4. Disconnect RP1 hose at the plastic disconnect attached to the slow-start throttle valve.
  3140   Tue Jun 29 23:49:18 2010 steve, ranaUpdateVACslow pumpdown started

Untitled.png

The pumpdown started at 4 PM (2300 UTC). At 10 PM, we (Jenne, Jan, and I) opened up the RV1 valve to full open. That's the second inflection point in the plot.

  3139   Tue Jun 29 20:47:19 2010 JenneUpdatePEMTerminator put on Guralp Box

A little D-sub terminator was put on the Gur1 input to the Guralp box, to check again the noise level of the box.

  3138   Tue Jun 29 17:10:49 2010 steve, ranaUpdateVACslow pumpdown started

The folding crane was fixed and tested this morning by the NNN rigging company. Pictures will be posted by Steve in the morning.

Afterwards, the ITM-east door was installed, jam-nuts checked. No high voltage was on for the in-vac PZTs.

The annulus spaces were roughed down to 350mTorr by Roughing Pump RP1. For this operation, we removed the low flow valve from the RP1 line.

After the spaces came down to ~400 mTorr, we closed their individual valves.

Warning: The VABSSC1 and VABSSC0 valves are incorrect and misleadingly drawn on the Vacuum overview screen.

We then:

  1. Closed V6 (valve between RP1 and the annulus line).
  2. Turned off RP1 from the MEDM screen.
  3. Installed the soft -starting butterfly valve.
  4. Turned on RP1.
  5. Opened V3.
  6. Closed VV1 (at the last minute - this is a vent valve and must be checked before each pumpdown)
  7. and pumpdown was started with a 3/4 turn opening of manual valve RV1.

Our idea is to have a much slower pumpdown this time than the last time when we had a hurricane kick up the dust. Looks like it worked, but next time we should do only 1/2 turn.

  3137   Tue Jun 29 16:44:12 2010 Jenne, ranaUpdateMOPAMOPA is NOT dead, was just asleep

Quote:

Not dead. It just had a HT fault. You can tell by reading the front panel. Cycling the power usually fixes this.

MOPA is back onliine.  Rana found that the fuse in the AC power connector's fuse had blown.  This was evident by smelling all of the inputs and outputs of the MOPA controller. The power cord we were using for this was only rated for 10A and therefore was a safety hazard. The fuse should be rated to blow before the power cord catches on fire. The power cord end was slightly melted. I don't know why it hadn't failed in the last 12 years, but I guess the MOPA was drawing a lot of extra current for the DTEC or something due to the high temperature of the head.

We got some new fuses from Todd @ Downs. 

The ones we got however were fast-blow, and that's what we want  The fuses are 10A, 250V.  The fuses are ~.08 inches long, 0.2 inches in diameter. 

  3136   Tue Jun 29 14:19:44 2010 josephbUpdateCDSDaily Downs Update (Part 2)

I picked up the ribbon cable connectors from Jay.  It looks like we'll have to make the new cables for connecting the ADCs/DACs myself (or maybe with some help).  We should be able to make enough ribbon cables for use now.  However,  I'm adding "Make nice shielded cables" to my long term to do list.

I pointed out the 2 missing adapter boxes we need to Jay.  He has the parts (I saw them) and will try to get someone to put it together in the next day or so.  I also picked up 2 more D080281 (DB44 to SCSI), giving us enough of those.

I once again asked Jay for an update on IO chassis, and expressed concern that without them the CDS effort can't really go forward, and that we really need this to come together ASAP.  He said they still need to make 3 new ones for us.

So we're still waiting on a computer, 3 IO chassis, router + ethernet.

  3135   Tue Jun 29 14:16:35 2010 KojiConfigurationSAFETYBack in LASER HAZARD mode.

The insects and the laser trouble... Strange coincidences with LHO surprised me, but now I have been relieved.

Quote:

[Steve, Kiwamu, Jenne]

The 40m is now back in Laser Hazard mode.  Safety glasses are required for entry into the LVEA / IFO room.

 

  3134   Tue Jun 29 12:08:43 2010 josephbUpdateCDSDaily Downs Update

I talked with Rolf, and asked if we were using Megatron for IO.  The gist boiled down to we (the 40m) needed to use it for something, so yes, use it for the IO computer.  In regards to the other end station computer, he said he just needed a couple of days to make sure it doesn't have anything on it they need and to free it up.

I had a chat with Jay where he explained exactly what boards and cables we need.  Adapter boards are 95% of the way there.  I'll be stopping by this afternoon to collect the last few I need (my error this morning, not Jays).  However it looks like we're woefully short on cables and we'll have to make them. I also acquired 2 D080281 (Dsub 44 x2 to SCSI).

For each 2 Pentek DACs plus a 110B, you need 1 DAC adapter board (D080303 with 2 connectors for IDC40 and a SCSI).  You also need a D080281 to plug onto the back of the Sander box (going to the 110Bs) to convert the D-sub 44 pins to SCSI.

LSC will need none, SUS will need 3, IO will need 1, and the ends will need 1 each.  We have a total of 6, we're set on D080303s.  We have 3 110Bs, so we need one more D080281 (Dsub44 to SCSI).  I'll get that this afternoon.

For each XVME220, we'll need one D080478 binary adapter.  We have 8 XVME220s, and we have 8 boards, so we're set on D08478s.

For the ends, there's a special ADC to DB44/37 adapter, which we only have 1 one of.  I need to get them to make 1 more of these boxes.

We have 1 ADC to DB37 adapter, of which we'll need  1 more of as well, one for IO and one for SUS. 

However, for each Pentek ADC, we need a IDC40 to DB37 cable.  For each Pentek DAC we need an IDC40 to IDC40 cable.  We need a SCSI cable for each 110B.  I believe the current XVME220 cables plug directly in the BIO adapter boxes, so those are set.

So we need to make or acquire 11 IDC40 to DB37 cables, 7 IDC40 to IDC40 cables, and 3 SCSI cables.

 

Summary Needed:

1 ADC to DB44/37 for the End (D080397)

1 ADC adapter (D080302)

1 Dsub44 to SCSI (D080291)

11 IDC40 to DB37 cables

7 IDC40 to IDC40 calbes

3 SCSI cables

PLUS from before:

3 IO Chassis (2 Dolphin, 1 Small)

1 1U computer (8 core for end)

Router/2 50+m ethernet for DAQ

  3133   Tue Jun 29 11:48:17 2010 JenneConfigurationSAFETYBack in LASER HAZARD mode.

[Steve, Kiwamu, Jenne]

The 40m is now back in Laser Hazard mode.  Safety glasses are required for entry into the LVEA / IFO room.

  3132   Tue Jun 29 10:20:58 2010 ranaUpdateMOPAMOPA is NOT dead

Not dead. It just had a HT fault. You can tell by reading the front panel. Cycling the power usually fixes this.

  3131   Tue Jun 29 08:55:18 2010 JenneFrogsEnvironmentWe're being attacked!

Infested_InvasionOfKillerBugs.jpg

We're going to have to reinstate the policy of No food / organic trash *anywhere* in the 40m.  Everyone has been pretty good, keeping the food trash to the one can right next to the sink, but that is no longer sufficient, since we've been invaded by an army of ants:

AntInvasion_small.jpg

We are going back to the old policy of Take your trash out to the dumpsters outside.  I'm sure there are some old wives tales about how exercise after eating helps your digestion, or something like that, so no more laziness allowed!

  3130   Tue Jun 29 08:41:06 2010 steveUpdateMOPAMOPA is dead

I found the laser dead this morning.

The crane people are here to unjam it.

Laser hazard mode is lifted and LASER SAFE MODE is in place. No safety glasses but CRANE HAZARD is still active.

Stay out of the 40m lab !

 

 

Attachment 1: laserisdead.jpg
laserisdead.jpg
  3129   Mon Jun 28 21:26:05 2010 ranaSummaryCDSCDS adapter board notes

Those drawings are an OK start, but its obvious that things have changed at the 40m since 2002. We cannot rely on these drawings to determine all of the channel counts, etc.

I thought we had already been through all this...If not, we'll have to spend one afternoon going around and marking it all up. 

  3128   Mon Jun 28 13:40:53 2010 josephb, AlexUpdateCDSChanges added to CDS SVN, new checkout, new features, some changes made

Last week Alex merged in the changes I had made to the local 40m copy of the Real Time Code Generator.  These were to add a new part, called FiltMuxMatrix, which is a matrix of filter banks, as well as fixing the filter medm generation code so the filter banks actually have working time stamps.

I checked out a new version of the CDS SVN with these changes merged in.  Changes that will be added in the near future by Rolf and Alex include the addition of "tags".  These are pieces in simulink which act as a bridge between two points, so you can reduce the amount of wire clutter on diagrams.  Otherwise they have no real affect on the generated C code.  Also the ADC/DAC channel selector and in fact the ADC/DAC parts will be changing.  The MIT group has requested the channel selector be freed up for its original purpose in matlab, so Rolf is working on that.

The new checkout includes the new directory scheme Rolf is pushing.  So when you run the code generator and more specifically, install SYS, it places code in /opt/rtcds/caltech/c1/ type directories, like medm, chans, target, scripts.

For the time being, Alex has created a directory /rtcds on Linux1 under /home/cds.  He then created softlinks to that directory on megatron, c1iscex, and allegra in the /opt directory.  This was an easy way to have a shared path.

However, it does mean on each new FE  machine after setting up the mounting of /home/cds from Linux1, we also need to create the /opt/rtcds link to /cvs/cds/rtcds.

After checking out the CDS SVN, we discovered there some files missing that Alex had added to his version, but not the main branch.  Alex came over to the 40m and proceeded to get all those files checked in.  We then checked it out again.  Changes were made to the awg, framebuilder, and nds codes and needed to be rebuilt. 

There's a new naming scheme for models.  You need to include the site before the 3 letter system name.  So lsc.mdl become c1lsc.mdl.

Certain other file name conventions were also changed.  Instead of tpchn_c1.par, tpchn_c2.par, etc, its now tpchn_c1lsc.par, tpchn_c2lsp.par, etc.  The system name is included at the end of the filename, to help make it clearer what file goes with what.

This required an edit of the chnconf file, which has explicit calls to those file names.  Once we edited that file, we had to reload the xinetd service which its apparently a subpart of (this can be accomplished by /etc/init.d/xinetd stop, then /etc/init.d/xinetd start).

/etc/rc.d/rc.local also had to be edited for the new model names (c1lsc, c1lsp, etc).

The daqdrc file (for the framebuilder) now parses which dcu_rate to use from the tpchn_c1lsc.par type files, so the dcu_rate 20 = 16384 lines have been removed.  set gds_server has also been removed, and replaced with tpconfig "/opt/rtcds/caltech/c1/target/gds/param/testpoint.par";  from which it can get the hostname.  This information is now derived from the c1SYS.mdl file.

Hostname needs to be added to the .mdl files, in the cdsParameter block (i.e. host=megatron).

After that Alex informed me the IOP processor needs to be running for the other models to work properly, as well as for the Framebuilder to work.

The models and framebuilder now get their timing signal from the IOP (input/output processor).  This must be running in order for the other models or FB to run properly.  Its generally named c1x00 or c1x01 or similar.  The last two numbers ideally are unique to each FE computer.

Initially there was a problem running on Megatron, because the IOP gets its timing signal from the IO chassis, and there was none connected to megatron.  However, he has since modified the code so that if there's no IO chassis, the IOP processor just uses the system clock.  It has been tested and runs on megatron now.

 

  3127   Mon Jun 28 12:48:04 2010 josephbSummaryCDSCDS adapter board notes

The following is according to the drawing by Ben Abbott found at http://www.ligo.caltech.edu/~babbott/40m_sus_wiring.pdf

This applies to SUS:

Two ICS 110Bs.  Each has 2 (4 total) 44 shielded conductors going to DAQ Interface Chassis  (D990147-A).  See pages 2 and 4.

Three Pentek 6102 Analog Outputs to LSC Anti-Image Board (D000186 Rev A).  Each connected via 40 conductor ribbon cable (so 3 total). See page 5.

Eight XY220 to various whitening and dewhitening filters.  50 conductor ribbon cable for each (8 total). See page 10.

Three Pentek 6102 Analog Input to Op Lev interface board. 40 conductor ribbon cable for each (3 total).  See page 13.

 

The following look to be part of the AUX crate, and thus don't need replacement:

Five VMIC113A to various Coil Drives, Optical Levers, and Whitening boards.  64 conductor ribbon cable for each (5 total). See page 11.

Three XY220 to various Coil boards. 50 conductor ribbon for each (3 total).  See page 11.

The following is according to the drawing by Jay found at http://www.ligo.caltech.edu/~jay/drawings/d020006-03.pdf

This applies to WFS and LSC:

Two XY220 to whitening 1 and 2 boards.  50 conductor ribbon for each (2 total).  See page 3.

Pentek 6102 to LSC Anti-image. 50 conductor ribbon. (1 total). See page 5.

 

The following are unclear if they belong to the FE or the Aux crate.  Unable to check the physical setup at the moment.

One VMIC3113A to LSC I & Q, RFAM, QPD INT. 64 conductor ribbon cable. (Total 1).  See page 4.

One XY220 to QPD Int.  50 conductor ribbon cable. (Total 1). See page 4.

 

The following look to be part of WFS, and aren't needed:

Two Pentek 6102 Analog Input to WFS boards. 40 conductor ribbon cables (2 Total). See page 1.

The following are part of the Aux crate, and don't need to be replaced:

Two VMIC3113A to Demods, PD, MC servo amp, PZT driver, Anti-imaging board. 64 conductor ribbon cable (2 Total). See page 3.

Two XY220 to Demods, MC Servo Amp, QPD Int boards.  50 conductor ribbon cable (2 Total). See page 3.

Three VMIC4116 to Demod and whitening boards.  50 conductor ribbon cable (3 Total). See page 3.

  3126   Mon Jun 28 11:27:08 2010 MeganUpdateElectronicsMarconi Phase Noise

Using the three Marconis in 40m at 11.1 MHz, the Three Cornered Hat technique was used to find the individual noise of each Marconi with different offset ranges and the direct/indirect frequency source of the rubidium clock.

Rana explained the TCH technique earlier - by measuring the phase noise of each pair of Marconis, the individual phase noise can be calculated by:

S1 = sqrt( (S12^2 + S13^2 - S23^2) / 2)

S2 = sqrt( (S12^2 + S23^2 - S13^2) / 2)

S3 = sqrt( (S13^2 + S23^2 - S12^2) / 2)

I measured the phase noise for offset ranges of 1Hz, 10Hz, 1kHz, and 100kHz (the maximum allowed for a frequency of 11.1Mhz) and calculated the individual phase noise for each source (using 7 averages, which gives all the spikes in the individual noise curves). The noise from each source is very similar, although not quite identical, while the noise is greater at higher frequencies for higher offset ranges, so the lowest possible offset range should be used. It appears the noise below a range of 10Hz is fairly constant, with a smoother curve at 10Hz.

The phase noise for direct vs indirect frequency source was measured with an offset range of 10Hz. While very similar at high and low frequencies for all 3 Marconis, the indirect source was consistently noisier in the middle frequencies, indicating that any Marconis connected to the rubidium clock should use the rubidium clock as a direct frequency reference.

Since I can't adjust settings of the Marconis at the moment, I have yet to finish measurements of the phase noise at 160 MHz and 80 MHz (those used in the PSL lab), but using the data I have for only the first 2 Marconis (so I can't finish the TCH technique), the phase noise appears to be lowest using the 100kHz offset except at the higher frequencies. The 160 MHz signal so far is noisier than the 11.1 MHz signal with offset ranges of 1 kHz and 10 Hz, but less noisy with a 100 kHz offset.

I still haven't measured anything at 80 MHz and have to finish taking more data to be able to use the TCH technique at 160 MHz, then the individual phase noise data will be used to measure the noise of the function generators used in the PSL lab.

Attachment 1: IndividualNoise11100kHzAllRanges.jpg
IndividualNoise11100kHzAllRanges.jpg
Attachment 2: IndividualNoise11100kHzSeparate.jpg
IndividualNoise11100kHzSeparate.jpg
Attachment 3: DirectvsIndirectNoise.jpg
DirectvsIndirectNoise.jpg
Attachment 4: FG12Noise.jpg
FG12Noise.jpg
  3125   Sat Jun 26 21:13:19 2010 ranaSummaryComputer Scripts / ProgramsCOMSOL 4.0 Installation

I've installed COMSOL 4.0 for 32/64 bit Linux in /cvs/cds/caltech/apps/linux64/COMSOL40/

It seems to work, sort of.


Notes:

  1. It did NOT work according to the instructions. The CentOS automount had mounted /dev/scd0 on /media/COMSOL40. In this configuration, I was getting a permission denied error when trying to run the default setup script. I did a 'sudo umount /dev/scd0' to get rid of this bad mount and then remounted using 'sudo mount /dev/dvd /mnt'. After doing this, I ran the setup script '/mnt/setup' and got the GUI which started installing as usual.
  2. I also pointed it at the linux64/matlab/ installation.
  3. It seems to not work right on Rosalba because of my previous java episode. The x-forwarding from megatron also fails. It does work on allegra, however.
  3124   Sat Jun 26 20:16:44 2010 josephbUpdateCDSNew checkout of RCG from SVN and changes made

ORPHAN ENTRY FOUND ON ROSALBA:::::::::::::::::::::::::::::::::::::::::::::::::::>>>>>>>>>>>>>>>>>>>>>>>>>>>>>


We did svn update.  Then Alex realized he missed adding some files, so added them to the svn, and then we checked out again.

 

We rebuilt awg, fb, nds.

We reloaded service xinet.d, after editing /etc/xinetd.d/chnconf.  We changes all the tpchn_c1 to tpchn_c1SYS

 

There's a new naming scheme for the model files.  They start with the site name.  So, lsc becomes c1lsc. 

 

On any machine you want code running, make a symbolic link to /cvs/cds/rtcds/ in /opt/

  3123   Sat Jun 26 05:02:04 2010 ranaHowToSVNSVN woes

Quote:

I am trying to get an actual complete install of the 40m svn on my machine. It keeps stopping at the same point:

 I have always seen this when checking out the 40m medm SVN on a non-Linux box. I don't know what it is, but Yoichi and I investigated it at some point and couldn't reproduce it on CentOS. I think its some weirdness in the permissions of tmp files. It can probably be fixed by doing some clever checkin from the control room.

Even worse is that it looks like the whole 'SVN' mantra has been violated in the medm directory by the 'newCDS' team. It could be that Joe has decided to make the 40m a part of the official CDS SVN, which is OK, but will take some retraining on our part.

  3122   Fri Jun 25 20:32:30 2010 kiwamuUpdateGreen Lockinggreen power on the PSL table

The power of the green beam generated on the PSL table should be about 650uW in terms of the shot noise.

       One of the important parameters we should know is the power of the green beam on the PSL table because it determines the SNR.

The green beam finally goes to a photo detector together with another green beam coming from the arm cavity, and they make a beat signal and also shot noise.

So in order to obtain a good SNR toward the shot noise at the photo detector, we have to optimize the powers.

If we assume the green power from the arm is about 650uW,  a reasonable SNR can be achieved when these powers are at the same level. 

To get such power on the PSL table, a 90% partial reflector is needed for picking it off from the PSL as we expected.

 


  power dependency of SNR

      Suppose two lasers are going to a photo detector while they are beating (interfering).

The beat signal is roughly expressed by

      [signal]  ~ E1EE1 E2*,

                     ~ 2 ( PP2)½ cos (phi), 

 where  E1 and Erepresent the complex fileds,  Pand Prepresent their powers and phi is a phase difference.

This equation tells us that the strength of the signal is proportional to  ( PP2)½  .

At the same time we will also have the shot noise whose noise level depends on the inverse square route of the total power;

          [noise] ~ ( PP2)½.

 According to the equations above, SNR is expressed by

        SNR = [signal] / [noise] ~ ( PP2)½  / ( PP2)½.

If we assume Pis fixed,  the maximum SNR can be achieved when  P2 goes to the infinity. But this is practically impossible.

Now let's see how the SNR grows up as the power P increases. There are two kinds of the growing phase.

    (1) When PP1 , SNR is efficiently improved with the speed of  P2½.

    (2) But  when P>   P1 , the speed of growing up becomes very slow. In this regime increasing of  P2 is highly inefficient for improvement of the SNR.

Thus practically PP is a good condition for the SNR.

At this point the SNR already reaches about 0.7 times of the maximum, it's reasonably good.

 


 power estimation

         According to the fact above, we just adjust the green powers to have the same power levels on the PSL table.

 The table below shows some parameters I assume when calculating the powers.

ITM transmissivity @ 532nm  Ti 1.5 %
ETM transmissivity @ 532nm Te 4.5 %
Transmissivity of the arm cavity @ 532nm T_cav 74.4 %
Transmissivity of the BS @ 532nm T_BS 97 %
Transmissivity of  PR1 and SR1 @ 532nm T_PR 90%
Transmissivity of the PMC @ 1064 nm T_pmc 65 %
The power of the green beam at the end station P_end 1 mW
The power of the PSL  P_psl 2 W
Conversion efficiency of the PPKTP eta 3 %/W

         Attached figure shows a simplified schematic of the optical layout with some numbers. 

By using those parameters we can find that the green beam from the arm cavity is reduced to 650uW when it reaches the PSL table.

To create the green beam with the same power level on the table, the power of 1064 nm going to the doubling crystal should be about 150mW.

This amount of the power will be provided by putting a 90% partial reflector after the PMC.

 

Attachment 1: optical_power.png
optical_power.png
  3121   Fri Jun 25 15:22:45 2010 JenneOmnistructureSAFETYNo entry to the 40m LVEA until further notice!

                                                                                                                                                                                                                                                                               

The 40m corner station crane is out of order, and it's stuck in a way that prohibits entry to the 40m LVEA / IFO room for safety.  The crane has been locked out / tagged out. 

Until further notice, absolutely no one may enter the 40m LVEA.  Work is permitted in the desk / control room areas.

Signs have been posted on all doors into the LVEA.  Please consider those doors locked out / tagged out.

                                                                                                                                                                                                                                                                               

ELOG V3.1.3-